A wine company needs to blend a California wine with a 5% alcohol content and a French wine with a 9% alcohol content to obtain 200 gallons of wine with a 7% alcohol content. How many gallons of each kind of wine must be used?
Answer by jorel1380(3719)

200x.07=14
m+n=200
.05m+.09n=14
5m+9n=1400
5m+5n=1000
4n=400
n=100
m=100
You need 100 gallons of each kind to make the desired mixture..